# Weakly Supervised Gaussian Contrastive Grounding with Large Multimodal Models for Video Question Answering

### v1

@@ 4.2 Pseudo Labels for Temporal Grounding
Event Description Generation. To adapt the textual representation for better event description and reduce the semantic gaps, we prompt the LLaMA Touvron et al. (2023) (or any other open-sourced LLM) to fuse the question and answer pairs with hand-written demonstrations. For example, the QA-pair [Q: How does the boy react after opening the present? A: carry up the dog] will be transformed into the declarative event description [The boy carries up the dog after opening the present.]. The event descriptions provide more accurate textual descriptions for question-critical moments because the answer content is included.
@@
Pseudo Labels Generation. As in Figure 2 (b), we represent the video and event description with the class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ and $\textbf{d}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ respectively, to obtain the weakly labeled question-critical timestamps. In detail, we compute the cosine similarities between $\textbf{E}_{[CLS]}$ and $\textbf{d}_{[CLS]}$ and get the similarities scores $s\in\mathbb{R}^{T}$, which recording the relevance between each frame and the event description. Then, we choose the indexes of the highest Top-$K$ scores in $s$ as $w\in\mathbb{R}^{K}$ where $w_{k}\in\left\{1,2,\cdots,T\right\}$, to be the timestamps of question-critical frames. We verify the effectiveness of the weakly grounded $w$ in the analysis 5.4.
4.3 Gaussian Generator
@@
Recent research Xiao et al. (2023) highlights the superiority of end-to-end Gaussian mask learning for grounding in VideoQA. Motivated by this, we design the effective weakly supervised Gaussian-based Contrastive Grounding (GCG). However, unlike Xiao et al. (2023) to generate a single Gaussian function, we generate multiple Gaussian functions to characterize the multi-event temporal structure of the video. Moreover, our GCG is optimized from both the QA supervision $\mathcal{L}_{vqa}$ and weakly labeled supervision, which is more effective for discovering the question-critical moments.
@@
Specifically, we utilize the Gaussian generator to obtain $K$ Gaussian functions $g=\left\{g_{1},\cdots,g_{K}\right\},g_{k}\in\mathbb{R}^{T}$, depending on the video and question. These Gaussian functions will be combined into an overall weight distribution $p\in\mathbb{R}^{T}$ for each video moment. Notably, $p$ tends to have $K$ peaks, representing the most question-critical $K$ frames, with corresponding indexes to be the centers of each Gaussian function.
@@
As in the right part of Figure 3, the Gaussian generator consists of a cross-modal embedding layer and a transformer encoder Vaswani et al. (2017). The cross-modal embedding layer is a down-sampling linear layer with learnable modal-type embeddings and positional embeddings. The concatenated multimodal embeddings $\textbf{M}=[\textbf{E}_{[CLS]};\textbf{Q}]\in\mathbb{R}^{(T+L_{q})\times D_{I}}$ serve as inputs for the Gaussian generator:
M $=\operatorname{Linear}(\textbf{M}),\ \textbf{M}\in\mathbb{R}^{(T+L_{q})\times D_{G}}$ (2)
@@
$\mu=\operatorname{Sigmoid}(\operatorname{Linear}(\textbf{G})),\ \mu\in\mathbb{R}^{K}$ (3)
Given predicted $\mu$, we get $K$ Gaussian functions $g=\left\{g_{1},\cdots,g_{K}\right\}$ parameterized with ($\mu$, $\sigma$):
$g_{k}$ $=\frac{1}{\sqrt{2\pi}\sigma}\mathrm{exp}(-\frac{(t/T-\mu_{k})^{2}}{2\sigma^{2}}),\ g_{k}\in\mathbb{R}^{T}$ (4)
@@
$p=\mathrm{Norm}(\sum_{k=1}^{K}g_{k}),\ p\in\mathbb{R}^{T}$ (5)
$\mathrm{Norm}$(·) scales values into the range [0, 1]. As the $K$ peaks in $p$, whose corresponding indexes are $\left\{\mu_{1},\cdots,\mu_{K}\right\}$, represent the most question-critical $K$ frames, we optimize the Gaussian generator with the regression objective to measure the discrepancy between the predicted centers $\mu\in\mathbb{R}^{K}$ and the weakly grounded timestamps $w\in\mathbb{R}^{K}$ by smooth $L1$ loss:
$\mathcal{L}_{reg}=\sum_{k=1}^{K}\mathrm{Smooth}_{L1}\|\mu_{k}-w_{k}/T\|$ (6)
4.4 Contrastive Grounding
@@
Positive Moments Selection. Since the distribution $p$ weights each video moment based on its contribution to the question, we select the Top-$K$ elements from $\textbf{E}_{[CLS]}$ according to $p$, and obtain $\textbf{E}^{pos}\in\mathbb{R}^{K\times D_{I}}$ as the positive frames. However, the selection via vanilla hard Top-$K$ produces a discrete selection, making it inapplicable for end-to-end training. We address this issue by adopting a differentiable Top-$K$ using the perturbed maximum method Berthet et al. (2020).
Negative Moments Mining. To distinguish highly confusing scenes, we mine negative moments within the same video as intra-negative frames $\textbf{E}^{intra}\in\mathbb{R}^{N_{intra}\times D_{I}}$, by sampling frames with the lowest $N_{intra}$ weights in $p$ from $\textbf{E}_{[CLS]}$. We also use $N_{inter}$ frames randomly sampled from other videos within the same batch to serve as inter-negative frames $\textbf{E}^{inter}\in\mathbb{R}^{N_{intra}\times D_{I}}$. These negative samples from both the same video and other videos can provide richer information. The objective of contrastive grounding is described as:
$\mathcal{L}_{con}=-\frac{1}{K}\sum_{k=1}^{K}\mathrm{log}\frac{\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{pos}_{k}/\tau)}{\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{pos}_{k}/\tau)+\mathrm{SUM}}$ (7)
@@
where $\tau$ is the temperature factor and $\otimes$ is the dot product. Contrastive grounding can maximize the similarity between the query $\textbf{d}_{[CLS]}$ and a group of corresponding positive video moments $\textbf{E}^{pos}$ under the joint embedding space while pushing away negative video moments.
4.5 Answer Prediction
@@
$\mathcal{L}=\mathcal{L}_{vqa}+\alpha_{1}\mathcal{L}_{reg}+\alpha_{2}\mathcal{L}_{con}$ (8)
where $\alpha_{1}$ and $\alpha_{2}$ are the hyper-parameters to control the strengths of GCG. During the inference process, GCG only generates the weight distribution $p$ with the Gaussian generator as in Section 4.3, and obtains the most weighted $K$ frame embeddings $\hat{\textbf{E}}$ via a fully discrete Top-$K$ selection.
5 Experiments
5.1 Datasets
NExT-QA Xiao et al. (2021) contains 5.4k videos with an average length of 44s and 52k QA pairs, including question types of description, causal, and temporal. Intent-QA Li et al. (2023b) focuses on intent reasoning in daily social activities, with more than 4.3k videos and 16k QA pairs, including question types of causal-why, causal-how, and temporal. Causal-VidQA Li et al. (2022a) selects 27k video clips and asks 108k questions, including types of description, explanation, prediction, and counterfactual. MSVD-QA and MSRVTT-QA Xu et al. (2017) emphasize the description of video objects, activities, and their attributes, with 50k QA pairs over 1,970 videos and 243K QA pairs over 10K videos respectively. ActivityNet-QA Yu et al. (2019) consists of 58k QA pairs on 5.8k long web videos, with an average length of 180 seconds. NExT-QA, Intent-QA, and Causal-VidQA use a multi-choice setting to test temporal reasoning with causal and commonsense relations. MSVD-QA, MSRVTT-QA, and ActivityNet-QA employ an open-ended setting, focusing on the description of different elements in videos.
5.2 Implementation Details
We choose InstructBLIP Dai et al. (2023) as our LMM, with EVA-CLIP Fang et al. (2023) as the image encoder and text encoder, Q-former Li et al. (2023c) as the connection module, and Vicuna Zheng et al. (2023) or FLAN-T5 Chung et al. (2022) as the large language model. We sample each video as a sequence of $T=32$ frames and select $K=4$ frames as visual inputs. The number of negative samples is $N_{intra}=16$ and $N_{inter}=32$. The number of transformer encoder layers in the Gaussian generator is 2, with the hidden size $D_{G}=256$. For the hyperparameters, we set $\sigma=0.2$, $\tau=0.1$, $\alpha_{1}=\alpha_{2}=0.1$. During training, we keep the parameters of the image encoder, LLM, and text encoder frozen. We use AdamW to optimize the model with a learning rate of $1e^{-5}$ and the strategy of mixed precision.
Method NExT-QA Causal-VidQA Intent-QA
@@ 5.2 Implementation Details
Co-Mem Gao et al. (2018) 54.4 50.0 45.9 48.5 64.1 62.8 31.4 32.6 47.7 47.7 54.9 39.1 46.8
HCRN Le et al. (2020) 54.0 49.3 47.1 48.9 56.4 61.6 32.6 32.7 48.1 - - - -
HME Fan et al. (2019) 57.4 48.9 46.8 49.2 63.4 61.5 28.9 30.9 46.2 46.1 54.3 40.8 46.2
B2A Park et al. (2021) 58.3 49.0 47.4 49.6 66.2 62.9 31.2 35.2 49.1 - - - -
HGA Jiang and Han (2020) 57.8 49.1 48.1 50.0 65.7 63.5 32.2 34.3 48.9 44.9 51.0 39.6 44.6
IGV Li et al. (2022c) 59.6 51.7 48.6 51.3 65.9 62.1 35.0 31.2 48.6 - - - -
HQGA Xiao et al. (2022a) 59.4 52.3 49.0 51.8 - - - - - 48.2 54.3 41.7 47.7
MCR Zang et al. (2023) 62.3 52.0 49.2 52.4 67.5 65.6 37.8 33.4 51.1 - - - -
VGT Xiao et al. (2022b) 67.3 54.5 52.8 55.7 70.8 70.3 38.4 42.0 55.4 51.4 56.0 47.6 51.3
CaVIR Li et al. (2023b) - - - - - - - - - 58.4 65.5 50.5 57.6
TranSTR Li et al. (2023e) 70.0 60.2 59.7 61.5 73.6 75.8 48.9 50.3 62.2 - - - -
SeViLA Yu et al. (2023) - - - 71.5 - - - - - - - - -
InstructBLIP Dai et al. (2023) 79.8 70.5 71.5 72.5 79.5 81.4 64.7 56.8 70.6 73.0 70.2 68.8 71.5
GCG (Ours) 80.7 72.6 74.2 74.6 80.7 82.3 66.5 59.1 72.1 75.0 71.9 69.2 73.1
@@
Table 1: Accuracy (%) on NExT-QA, Causal-VidQA, and Intent-QA. Des, Tem, and Cau denote question types of Descriptive, Temporal, and Causal in NExT-QA. Des, Exp, Pre, and Cou denote question types of Description, Explanation, Prediction, and Counterfactual in Causal-VidQA. CW, CH, and Tem denote question types of Causal Why, Causal How, and Temporal in Intent-QA.
Method MSVD MSRVTT ActivityNet
MHN Peng et al. (2022) 40.4 38.6 -
VQA-T Yang et al. (2021) 46.3 41.5 38.9
TG-VQA Li et al. (2023a) 52.5 46.3 48.3
MuLTI-L Xu et al. (2023b) 54.7 47.8 -
FrozenBiLM Yang et al. (2022) 54.8 47.0 43.2
UMT-L Li et al. (2023d) 55.2 47.1 47.9
InternVideo Wang et al. (2022) 55.5 47.1 -
HiTea Ye et al. (2022) 55.6 45.9 46.4
mPLUG2 Xu et al. (2023a) 58.1 48.0 -
COSA-L Chen et al. (2023b) 58.6 48.8 49.2
VALOR-L Chen et al. (2023a) 60.0 49.2 48.6
InstructBLIP Dai et al. (2023) 59.7 47.0 46.3
GCG (Ours) 61.7 49.5 49.9
@@
Table 2: Accuracy (%) on open-ended VideoQA datasets.
Settings NExT-QA Intent-QA MSVD-QA MSRVTT-QA
Baseline 72.5 71.5 59.7 47.0
Pseudo-labeled (Q) 72.3(-0.2) 70.3(-1.2) 59.9(+0.2) 47.5(+0.5)
Pseudo-labeled (E) 79.1(+6.6) 78.3(+6.8) 64.3(+4.6) 51.8(+4.8)
Table 3: As a preliminary step, we analyze the performance upper bound with weakly labeled keyframes as visual inputs.
5.3 Results and Analysis

### v3

@@ 4.2. Pseudo Labels for Temporal Grounding
Event Description Generation. To adapt the textual representation for better event description and reduce the semantic gaps, we directly prompt the LLM inside the LMM (e.g., Vciuna (Zheng et al., 2023)) to fuse the question and answer pairs with hand-written demonstrations. For example, the QA-pair [Q: How does the boy react after opening the present? A: carry up the dog] will be transformed into the declarative event description [The boy carries up the dog after opening the present.]. Since the event description is composed of simple changes in the grammatical structure of the question $\mathcal{Q}$ and answer $\mathcal{A}$, most open-sourced or API-based LLMs can easily achieve this. Notably, the event descriptions provide more accurate textual descriptions for question-critical moments because the answer content is included.
@@
Pseudo Labels Generation. As in Figure 2 (b), we represent the video and event description with the class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ and $\textbf{d}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ respectively, to obtain the weakly labeled question-critical timestamps. In detail, we compute the cosine similarities between $\textbf{E}_{[CLS]}$ and $\textbf{d}_{[CLS]}$ and get the similaruity scores $s\in\mathbb{R}^{T}$, which recording the relevance between each frame and the event description. Then, we choose the indexes of the highest Top-$K$ scores in $s$ as $w\in\mathbb{R}^{K}$, where each element $w_{k}\in\left\{1,2,\cdots,T\right\}$, to be the timestamps of question-critical frames. We verify the effectiveness of the weakly grounded $w$ in the analysis 5.3.
4.3. Gaussian Generator
@@
Recent research (Xiao et al., 2023) highlights the superiority of end-to-end Gaussian mask learning for grounding in VideoQA. Motivated by this, we design the effective weakly supervised Gaussian-based Contrastive Grounding (GCG). However, unlike (Xiao et al., 2023) to generate a single Gaussian mask, we generate multiple Gaussian masks to characterize the multi-event temporal structure of the video. Moreover, our GCG is optimized from both the QA supervision $\mathcal{L}_{vqa}$ and weakly labeled supervision $\mathcal{L}_{reg}$ and $\mathcal{L}_{con}$, which is more effective for discovering the question-critical moments.
@@
Specifically, we utilize the Gaussian generator to obtain $K$ Gaussian masks $g=\left\{g_{1},\cdots,g_{K}\right\},g_{k}\in\mathbb{R}^{T}$, depending on the video and question. These Gaussian masks will be combined into an overall weight distribution $p\in\mathbb{R}^{T}$, to indicate the importance of each video moment. Notably, $p$ tends to have $K$ peaks, representing the most question-critical $K$ frames, with corresponding indexes to be the centers of each Gaussian function $g_{k}$.
@@
As in the right part of Figure 3, the Gaussian generator consists of a cross-modal embedding layer and a transformer encoder (Vaswani et al., 2017). The cross-modal embedding layer is a down-sampling linear layer with learnable modal-type embeddings and positional embeddings. The concatenated multimodal embeddings $\textbf{M}=[\textbf{E}_{[CLS]};\textbf{Q}]\in\mathbb{R}^{(T+L_{q})\times D_{I}}$ serve as inputs for the Gaussian generator:
(2) M $=\operatorname{Linear}(\textbf{M}),\ \textbf{M}\in\mathbb{R}^{(T+L_{q})\times D_{G}}$
@@
(3) $\mu=\operatorname{Sigmoid}(\operatorname{Linear}(\textbf{G})),\ \mu\in\mathbb{R}^{K}$
Given predicted $\mu$, we get $K$ Gaussian functions $g=\left\{g_{1},\cdots,g_{K}\right\}$ as masks, parameterized with ($\mu$, $\sigma$):
(4) $g_{k}$ $=\frac{1}{\sqrt{2\pi}\sigma}\mathrm{exp}(-\frac{(t/T-\mu_{k})^{2}}{2\sigma^{2}}),\ g_{k}\in\mathbb{R}^{T}$
@@
(5) $p=\mathrm{Norm}(\sum_{k=1}^{K}g_{k}),\ p\in\mathbb{R}^{T}$
$\mathrm{Norm}$(·) scales values into the range [0, 1]. As the $K$ peaks in $p$, whose corresponding indexes tend to be $\left\{\mu_{1},\cdots,\mu_{K}\right\}$, represent the most question-critical $K$ frames, we optimize the Gaussian generator with the regression objective to measure the discrepancy between the predicted centers $\mu\in\mathbb{R}^{K}$ and the weakly grounded timestamps $w\in\mathbb{R}^{K}$ by smooth $L1$ loss:
(6) $\mathcal{L}_{reg}=\sum_{k=1}^{K}\mathrm{Smooth}_{L1}\|\mu_{k}-w_{k}/T\|$
4.4. Contrastive Grounding
@@
Positive Moments Selection. Since the distribution $p$ weights each video moment based on its contribution to the question, we select the Top-$K$ elements from $\textbf{E}_{[CLS]}$ according to $p$, and obtain $\textbf{E}^{pos}\in\mathbb{R}^{K\times D_{I}}$ as the positive frames. However, the selection via vanilla hard Top-$K$ produces a discrete selection, making it inapplicable for end-to-end training. We address this issue by adopting a differentiable Top-$K$ using the perturbed maximum method (Berthet et al., 2020).
Negative Moments Mining. To distinguish highly confusing scenes, we mine negative moments within the same video as intra-negative frames $\textbf{E}^{intra}\in\mathbb{R}^{N_{intra}\times D_{I}}$, by sampling frames with the lowest $N_{intra}$ weights in $p$ from $\textbf{E}_{[CLS]}$. We also use $N_{inter}$ frames randomly sampled from other videos within the same batch to serve as inter-negative frames $\textbf{E}^{inter}\in\mathbb{R}^{N_{intra}\times D_{I}}$. These negative samples from both the same video and other videos can provide richer information. The objective is described as an infoNCE loss:
(7) $\mathcal{L}_{con}=-\frac{1}{K}\sum_{k=1}^{K}\mathrm{log}\frac{\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{pos}_{k}/\tau)}{\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{pos}_{k}/\tau)+\mathrm{SUM}}$
@@
$\tau$ is the temperature factor and $\otimes$ is the dot product. Contrastive grounding can maximize the similarity between the query $\textbf{d}_{[CLS]}$ and a group of corresponding positive video moments $\textbf{E}^{pos}$ under the joint embedding space while pushing away negative ones.
4.5. Answer Prediction
@@
(8) $\mathcal{L}=\mathcal{L}_{vqa}+\alpha_{1}\mathcal{L}_{reg}+\alpha_{2}\mathcal{L}_{con}$
Figure 4. As a preliminary step, we analyze the performance upper bound with weakly labeled keyframes as visual inputs.
$\alpha_{1}$ and $\alpha_{2}$ are the hyper-parameters to control the strengths of GCG. During the inference process, GCG only generates the distribution $p$ as in Section 4.3, and obtains the most weighted $K$ frame embeddings $\hat{\textbf{E}}$ via a fully discrete Top-$K$ selection.
5. Experiments
5.1. Datasets
NExT-QA (Xiao et al., 2021) contains 5.4k videos with an average length of 44s and 52k QA pairs, including question types of description, causal, and temporal. Intent-QA (Li et al., 2023d) focuses on intent reasoning in daily social activities, with more than 4.3k videos and 16k QA pairs, including question types of causal-why, causal-how, and temporal. Causal-VidQA (Li et al., 2022b) selects 27k video clips and asks 108k questions, including types of description, explanation, prediction, and counterfactual. MSVD-QA and MSRVTT-QA (Xu et al., 2017) emphasize the description of video objects, activities, and their attributes, with 50k QA pairs over 1,970 videos and 243K QA pairs over 10K videos respectively. ActivityNet-QA (Yu et al., 2019) consists of 58k QA pairs on 5.8k long web videos, with an average length of 180 seconds. NExT-QA, Intent-QA, and Causal-VidQA use a multi-choice setting to test temporal reasoning with causal and commonsense relations. MSVD-QA, MSRVTT-QA, and ActivityNet-QA employ an open-ended setting, focusing on descriptive questions of different elements.
5.2. Implementation Details
We choose InstructBLIP (Dai et al., 2023) and BLIP-2 (Li et al., 2023b) as our LMM for their representative structure and widespread use, with EVA-CLIP (Fang et al., 2023) as the image encoder, Q-former (Li et al., 2023b) as the connection module, and Vicuna (Zheng et al., 2023) or FLAN-T5 (Chung et al., 2022) as the large language model. Following previous works (Yu et al., 2023; Li et al., 2023e, b; Dai et al., 2023), we sample each video as a sequence of $T=32$ frames and select $K=4$ frames as visual inputs. The number of negative samples is $N_{intra}=16$ and $N_{inter}=32$. The number of transformer encoder layers in the Gaussian generator is 2, with the hidden size $D_{G}=256$. For the hyperparameters, we set $\sigma=0.2$, $\tau=0.1$, $\alpha_{1}=\alpha_{2}=0.1$. During training, we keep the parameters of the image encoder, LLM, and text encoder frozen. We use AdamW to optimize the model with a learning rate of $1e^{-5}$ and the strategy of mixed precision. For multi-choice datasets like NExT-QA, Causal-VidQA, and Intent-QA, we use BLIP2-FLAN-T5-XL and IntructBLIP-FLAN-T5-XL. For open-ended datasets like MSVD-QA, MSRVTT-QA, and ActivityNet-QA, we use IntructBLIP-Vicuna-7B.
Method Source NExT-QA Causal-VidQA Intent-QA
@@ 5.2. Implementation Details
Co-Mem (Gao et al., 2018) CVPR’18 54.4 50.0 45.9 48.5 64.1 62.8 31.4 32.6 47.7 47.7 54.9 39.1 46.8
HCRN (Le et al., 2020) CVPR’20 54.0 49.3 47.1 48.9 56.4 61.6 32.6 32.7 48.1 - - - -
HGA (Jiang and Han, 2020) AAAI’20 57.8 49.1 48.1 50.0 65.7 63.5 32.2 34.3 48.9 44.9 51.0 39.6 44.6
IGV (Li et al., 2022d) CVPR’22 59.6 51.7 48.6 51.3 65.9 62.1 35.0 31.2 48.6 - - - -
HQGA (Xiao et al., 2022a) AAAI’22 59.4 52.3 49.0 51.8 - - - - - 48.2 54.3 41.7 47.7
B2A (Park et al., 2021) CVPR’21 58.3 49.0 47.4 49.6 66.2 62.9 31.2 35.2 49.1 - - - -
VCSR (Wei et al., 2023) ACMMM’23 62.3 51.5 53.0 54.1 66.0 65.4 41.2 34.1 51.7 - - - -
VGT (Xiao et al., 2022b) ECCV’22 67.3 54.5 52.8 55.7 70.8 70.3 38.4 42.0 55.4 51.4 56.0 47.6 51.3
CaVIR (Li et al., 2023d) ICCV’23 - - - - - - - - - 58.4 65.5 50.5 57.6
Raformer (Li et al., 2023f) ACMMM’23 67.8 57.7 58.2 59.6 71.8 73.8 41.2 48.9 58.9 - - - -
TranSTR (Li et al., 2023e) ICCV’23 70.0 60.2 59.7 61.5 73.6 75.8 48.9 50.3 62.2 - - - -
SeViLA (Yu et al., 2023) NIPS’23 80.8 66.4 71.9 71.5 - - - - - - - - -
BLIP-2 (Li et al., 2023b) ICML’23 79.4 64.9 69.7 69.6 78.4 80.9 65.1 56.4 70.1 74.2 67.1 66.0 71.0
+ GCG Ours 79.5 71.6 73.0 73.6 78.7 81.2 65.9 58.4 71.1 75.5 69.1 66.9 72.3
InstructBLIP (Dai et al., 2023) NIPS’23 79.8 70.5 71.5 72.5 79.5 81.4 64.7 56.8 70.6 73.0 70.3 68.8 71.5
+ GCG Ours 80.7 72.6 74.2 74.6 80.7 82.3 66.5 59.1 72.1 75.0 71.9 69.2 73.1
@@
Table 1. Accuracy (%) on NExT-QA, Causal-VidQA, and Intent-QA. Des, Tem, and Cau denote question types of Descriptive, Temporal, and Causal in NExT-QA. Des, Exp, Pre, and Cou denote question types of Description, Explanation, Prediction, and Counterfactual in Causal-VidQA. CW, CH, and Tem denote question types of Causal Why, Causal How, and Temporal in Intent-QA. We highlight the best results and second best results.
Method Source MSVD MSRVTT A-Net
PGAT (Peng et al., 2021) ACMMM’21 39.0 38.1 -
MHN (Peng et al., 2022) IJCAI’22 40.4 38.6 -
EIGV (Li et al., 2022c) ACMMM’22 43.7 39.3 -
Raformer (Li et al., 2023f) ACMMM’23 46.0 42.3 -
VQA-T (Yang et al., 2021) ICCV’21 46.3 41.5 38.9
TG-VQA (Li et al., 2023a) IJCAI’23 52.5 46.3 48.3
MuLTI-L (Xu et al., 2024) AAAI’24 54.7 47.8 -
FrozenBiLM (Yang et al., 2022) NIPS’22 54.8 47.0 43.2
UMT-L (Li et al., 2023c) ICCV’23 55.2 47.1 47.9
HiTea (Ye et al., 2023b) ICCV’23 55.6 45.9 46.4
mPLUG2 (Xu et al., 2023) ICML’23 58.1 48.0 -
COSA-L (Chen et al., 2024) ICLR’24 58.6 48.8 49.2
VALOR-L (Chen et al., 2023) Arxiv’23 60.0 49.2 48.6
InstructBLIP (Dai et al., 2023) NIPS’23 59.7 47.0 46.3
+ GCG Ours 61.7 49.5 49.9
@@
Table 2. Accuracy (%) on open-ended VideoQA datasets including MSVD-QA, MSRVTT-QA and ActivityNet-QA.
5.3. Pseudo-labels Analysis.
We first explore the performance of InstructBLIP with different frames as visual inputs to verify the effectiveness of the pseudo-labeled $w$: Pseudo-labeled (E) means we choose the $K$ frames whose indexes correspond to $w\in\mathbb{R}^{K}$ as visual inputs (detailed in Section 4.2). Pseudo-labeled (Q) is obtained the same way as Pseudo-labeled (E) but with the pure question for similarity computation. Baseline means we uniformly sample $K$ frames as visual inputs.
Figure 4 shows that Pseudo-labeled (E) exhibits significantly improved performance, particularly in benchmarks featuring longer videos and more complicated questions (+6.6% for NExT-QA and +6.8% for Intent-QA). This performance gap emphasizes the need and potential for more future work to effectively localize question-critical frames as visual inputs when using LMMs in video-language tasks. This also verifies the effectiveness of using the event description to provide pseudo-labels as weak supervision. Moreover, Pseudo-labeled (E) performs much better than Pseudo-labeled (Q). This can be explained from two perspectives: (1) The event descriptions include the contents of the answers needed to be grounded, filling the semantic gap between the pure question and the answer. (2) The CLIP models are mostly pre-trained on images and declarative texts, therefore the declarative event descriptions are more suitable for similarity computation to decide keyframes.
5.4. Main Results